OpenStudy (anonymous):

Russell runs 9/10 mile in five minutes. How many miles does he run in one minute

OpenStudy (alexandervonhumboldt2):

9/10 / 5=9/10*1/5=0.18 miles just divide the number of miles per 5 mins by number of mins

OpenStudy (anonymous):

Thank you sooo Much!! :)